Let f(x)=x^2-1 and g(x)=2x-3
Find (g o f)(3)
A. 0

B. 24

C. 13

D. 8

2.
Let f(x)=x^2 -1 and g(x)=2x+4
Find (f o g)(x)
A. 4x + 12

B. 4x2 - 12x + 4

C. 4x -12

D. 4x2 + 16x + 15


Answer by stanbon(75887)

Let f(x)=x^2-1 and g(x)=2x-3
Find (g o f)(3) = g[f(3)] = g[3^2-1] = g[8] = 2*8-3 = 13
-----------------
Ans: C
====================
A. 0
B. 24
C. 13
D. 8
2.
Let f(x)=x^2 -1 and g(x)=2x+4
Find (f o g)(x) = f[g(x)] = f[2x+4] = (2x+4)^2-1 = 4x^2+16x+15
===================================
Ans: D
================
Cheers,
Stan H.
=======================================
A. 4x + 12
B. 4x2 - 12x + 4
C. 4x -12
D. 4x2 + 16x + 15